Session: Feb 2011 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Pablo de Vicente / Paco Colomer at Tue Mar 01 15:01:58 2011) Success:

Medicina (Andrea Orlati at Tue Mar 01 15:11:29 2011) Minor failures:
vc02 locking problems

Shanghai (Bo Xia at Wed Mar 02 05:01:28 2011) Success:

Onsala (Michael Lindqvist at Wed Mar 02 05:25:11 2011) Success:
No known problems.

Hartebeesthoek (Jonathan Quick at Wed Mar 02 07:09:30 2011) Minor failures:RECE
Some gain instability observed in LCP channel. No other known problems

Torun (Kaz Borkowski at Wed Mar 02 07:43:05 2011) Success:
Like in previous NME experiment, our Mark5A wouldn't extract scan data for ftp-ing, but this time (thanks to Stephanie's recipe) we succeeded in manual extraction using DirList, play, data_check and disk2file commands (in this order) working in gaps between scans.

Effelsberg (Uwe Bach at Wed Mar 02 14:31:15 2011) Success:

Westerbork (Geert Kuper at Wed Mar 02 21:17:54 2011) Success:
No known problems

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:29:21 2011) Success:
Receiver may have been warm. No other known problems.